An inlet pump can fill a reservoir in 10 hours. An outlet pump can drain the same reservoir in 30 hours. How long would it take to fill the reservoir if both pumps were in use at the same time? Thank You
Answer by mananth(16949)

inlet pump fills in 10 hours
so it does 1/10 of the job in 1 hour.
outlet pump takes 30 hours
so it can do 1/30 of the job in 1 hour.
..
when both are open they will do 1/10 -1/30 of the job in 1 hour
=2/30 = 1/15 of the job in one hour.
so the tank will fill in 15 hours when both are on.
....
m.ananth@hotmail.ca